OpenStudy (anonymous):

Which of the following equations represents a line perpendicular to -6x + 7y = -1 ? Question 5 options: -6x - 35y = -1 -35x - 30y = -1 30x + 35y = -1 7x - 6y = -1

OpenStudy (anonymous):

First change the equation to y= form -6x+7y=-1 7y=6x-1 y=(6/7)x-(1/7) slope is 6/7 a line perpendicular to this has a slope that is the negative reciprocal of this so the slope will be -7/6 and if you change all of the choices to y= form, you'll find that -35x+30y=-1 has a slope of -7/6